# Relativistic and wide-angle corrections to galaxy power spectra

### v1

@@ 1 Introduction
A common approach to derive the wide-angle corrections to the Fourier power spectrum is to perturbatively compute the expansions of appropriate wide-angle terms in configuration space and then take the Fourier transform of the result [15, 16, 19, 23]. In this paper, we follow an alternative perturbative approach adopted by [22]. We generalise their result from the midpoint line of sight to all possible lines of sight. We include the standard redshift-space distortions and the non-integrated relativistic corrections from Doppler and Sachs-Wolfe effects, thereby generalising [23] which only considered the Doppler contribution. The integrated relativistic corrections from lensing magnification and time delay effects are included in the midpoint results of [22] but we omit these effects in our generalisation, leaving them for future work. We investigate the detectability of the relativistic, wide-angle and combined corrections by computing the $\chi^{2}$ for two futuristic galaxy surveys and their cross power.
@@ 2 Relativistic and wide-angle corrections in the number counts
For a given source population $a$, the dominant terms in the redshift-space number density contrast at linear order are
@@
where $|\bm{x}_{a}|$ is the comoving line-of-sight distance to redshift $z$, $b_{a}$ is the bias of tracer $a$, $\mathcal{H}=(1+z)^{-1}H=H_{0}\,[(1+z)^{3}\,\Omega_{m0}+1-\Omega_{m0}]^{1/2}$ is the conformal Hubble parameter, $\delta$ is the matter density contrast, $v_{i}=\partial_{i}V$ is the peculiar velocity, and $\partial_{\|}=\hat{\bm{x}}_{a}\cdot\bm{\nabla}$ is the derivative along the line of sight $\hat{\bm{x}}_{a}$. In Fourier space (2.1) is
@@
From now on, we will neglect terms that are of order $\delta(\mathcal{H}/k)^{n}$, where $n>2$. For consistency, all terms with $n\leq 2$ should be included.
@@
We follow [22] in neglecting the radial variations in a redshift bin, effectively treating all quantities in a bin as at the mid-redshift. See [23] for the inclusion of the radial corrections.
@@
$\frac{1}{x_{b}}$ $=$ $\frac{1}{r}\Big{[}1-t\big{(}\bm{\epsilon}\cdot\hat{\bm{r}}\big{)}+\frac{3}{2}t^{2}\big{(}\bm{\epsilon}\cdot\hat{\bm{r}}\big{)}^{2}-\frac{1}{2}t^{2}\epsilon^{2}\Big{]}\;,$ (2.15)
where we keep terms up to $O({\epsilon}^{2})$. Using the above results in (2.13), we obtain
@@
Then the full kernel, i.e. standard + combined relativistic wide-angle correction is
@@ 3 Auto- and cross-power spectra
where $P$ is the linear matter power spectrum and we omit the dependencies on $z$ here and below for brevity. These are the ‘local’ power spectra, i.e. the Fourier power spectra at line-of-sight position $\bm{r}(z)$. The ‘full’ power spectra are then given by an integral over all lines of sight [22],
@@
$\mathcal{K}_{a}(\bm{k}^{\prime},\bm{\epsilon},{\bm{r}})\,\mathcal{K}_{b}^{*}(\bm{k}^{\prime},\bm{\epsilon},{\bm{r}})=\sum_{l,m,n}\,\mathcal{C}_{lmn}^{(ab)}(\bm{k}^{\prime},{\bm{r}})\,\epsilon_{x}^{l}\,\epsilon_{y}^{m}\,\epsilon_{z}^{n}\;,$ (3.3)
where the coefficients $\mathcal{C}_{lmn}^{(ab)}$ are independent of $\bm{\epsilon}$ and where $l+m+n\leq 2$, since we expand up to $\epsilon^{2}$. Then the power spectrum (3.1) becomes
@@
$=P^{\mathrm{S}}_{ab}+P^{\mathrm{R}}_{ab}+P^{\mathrm{W}}_{ab}+P^{\mathrm{RW}}_{ab}+O\big{[}\epsilon^{3},\delta({\cal H}/k)^{3}\big{]}\;.$ (3.10)
@@
Detailed expressions for $\ell=0,1,2$ are given in Appendix C.
@@ 5 Detecting the relativistic and wide-angle effects
Our aim is to test whether the detection of relativistic and wide-angle effects is feasible in futuristic galaxy surveys. Following [26], we compute the $\chi^{2}$ values between the model $P_{ab}^{\rm S}+P_{ab}^{\rm I}$, which includes the relativistic and/or wide-angle effect I, and the standard model $P_{ab}^{\rm S}$, which excludes the effect I. We consider all the cases of $P_{ab}^{\rm I}$, where I=W, R, RW or W+R+RW. Our aim is to estimate the statistical significance of measuring effect I, against the null hypothesis of the absence of such an effect in the data. In our case, we use synthetic data generated by taking the theoretical prediction of the full power spectrum (3.10). Therefore the $\chi^{2}$ for effect I in a redshift bin centred at $z_{i}$ is simply
@@
$\mathrm{Var}\big{[}P_{aa}\big{]}=\frac{2}{N_{\bm{k}}}\,\big{|}\tilde{P}_{aa}\big{|}^{2}\,,\quad\mathrm{Var}\big{[}P_{12}\big{]}=\frac{1}{N_{\bm{k}}}\,\left[\,\big{|}\tilde{P}_{12}\big{|}^{2}+\big{|}\tilde{P}_{11}\big{|}\big{|}\tilde{P}_{22}\big{|}\,\right]\;,$ (5.2)
with
$\tilde{P}_{ab}=P_{ab}+\frac{\delta^{\mathrm{Kronecker}}_{ab}}{n_{a}}\qquad\mbox{and}\qquad P_{ab}=P_{ab}^{\rm S}+P_{ab}^{\rm R+W+RW}\,.$ (5.3)
@@
Note that $P_{ab}$ in the variance in (5.1)–(5.2) is given by (3.10) – i.e. it includes all relativistic and wide-angle effects. The number of Fourier modes is
$N_{\bm{k}}=\frac{2\,\pi\,k^{2}\,\Delta k\,\Delta\mu}{k_{\mathrm{f}}^{3}}\;,$ (5.4)
where the fundamental mode, $k_{\mathrm{f}}$, corresponds to the longest wavelength that can be measured in the redshift bin. The $\chi^{2}$ cumulative over redshift bins is
@@ Acknowledgements
SJ is supported by the Stellenbosch University Astrophysics Research Group fund. SG and RM are supported by the South African Radio Astronomy Observatory and National Research Foundation (grant no. 75415). SC acknowledges support from the Italian Ministry of University and Research, PRIN 2022 ‘EXSKALIBUR – Euclid-Cross-SKA: Likelihood Inference Building for Universe Research’, from the Italian Ministry of Foreign Affairs and International Cooperation (grant no. ZA23GR03), and from the European Union – Next Generation EU.
@@ Appendix A Fourier kernels
${}+2(t-\sigma_{a})\Big{[}k_{x}k_{z}\epsilon_{x}+k_{y}k_{z}\epsilon_{y}\Big{]}\Big{\}}+\mathcal{O}({\epsilon}^{3}),$
@@
${}+\frac{3}{2}\Omega_{m}\frac{\mathcal{H}^{2}}{k^{2}}\frac{(1-\mathcal{Q}_{a})}{r\mathcal{H}}\Big{[}2(t-\sigma_{a})\epsilon_{z}+(t-\sigma_{a})^{2}(\epsilon_{x}^{2}+\epsilon_{y}^{2}-2\epsilon_{z}^{2})\Big{]}+\mathcal{O}({\bm{\epsilon}}^{3}),\qquad$

### v2

@@ 1 Introduction
In this paper, we follow the perturbative approach adopted by [22]. We generalise their result from the midpoint line of sight to all possible lines of sight. We include the standard redshift-space distortions and the non-integrated relativistic corrections from Doppler and Sachs-Wolfe effects, thereby generalising [23] which only considered the Doppler contribution. The integrated relativistic corrections from lensing magnification and time delay effects are included in the midpoint results of [22] but we omit these effects in our generalisation, leaving them for future work. We investigate the detectability of the relativistic, wide-angle and combined corrections by computing the $\chi^{2}$ for two futuristic galaxy surveys and their cross power. If these corrections to the standard power spectra are detectable, then constraints on cosmological parameters could be affected when neglecting the corrections. In particular, we expect that measurements of the local primordial non-Gaussianity parameter, $f_{\rm NL}$, whose signal is strongest on very large scales, could be affected by relativistic wide-angle corrections.
We consider linear scalar perturbations on a late-time flat $\Lambda$CDM background, in the Newtonian gauge. The perturbed metric is
$\mathrm{d}s^{2}=a^{2}(\eta)\,\Big{[}-(1+2\Phi)\,\mathrm{d}\eta^{2}+(1-2\,\Phi)\,\mathrm{d}\bm{x}^{2}\Big{]},$ (1.1)
where $a$ is the scale factor, $\eta$ is conformal time and $\Phi$ is the gravitational potential. The matter 4-velocity is
$u^{\mu}=\frac{1}{a}\,\Big{(}1-\Phi,v^{i}\Big{)},$ (1.2)
where the peculiar velocity $v^{i}=\mathrm{d}x^{i}/\mathrm{d}\eta=a\,u^{i}$ is irrotational, so that $v_{i}=\partial_{i}V$, where $V$ is the velocity potential.
@@ 2 Relativistic and wide-angle corrections in the number counts
For a given source population $a$, the dominant terms in the redshift-space number density contrast at linear order (see [24] for beyond-linear effects) are
@@
where $|\bm{x}_{a}|$ is the background comoving line-of-sight distance to the background redshift $z$, $b_{a}$ is the bias of tracer $a$, $\mathcal{H}=(1+z)^{-1}\,H=H_{0}\,(1+z)^{-1}\,[(1+z)^{3}\,\Omega_{m0}+1-\Omega_{m0}]^{1/2}$ is the conformal Hubble parameter, $\delta$ is the comoving matter density contrast, and $\partial_{\|}=\hat{\bm{x}}_{a}\cdot\bm{\nabla}$ is the derivative along the line of sight $\hat{\bm{x}}_{a}$. In Fourier space (2.1) is
@@
From now on, we neglect terms in the 2-point correlations $\langle\Delta_{a}\,\Delta_{b}\rangle$, i.e. quadratic in the density contrast, that are of order $\delta(\mathcal{H}/k)^{n}$, where $n>2$. For consistency, all terms with $n\leq 2$ should be included in the 2-point correlations and power spectra.
@@
We follow [22] in neglecting the radial variations in a redshift bin, effectively treating all quantities in a bin as at the mid-redshift. Although it is straightforward to include these corrections (see [23]) they significantly complicate the expressions and we leave their inclusion for future work. An example of the effect of radial corrections is shown in Figure 5 below.
@@
$\frac{1}{x_{b}}$ $=$ $\frac{1}{r}\Big{[}1-t\big{(}\bm{\epsilon}\cdot\hat{\bm{r}}\big{)}+\frac{3}{2}t^{2}\big{(}\bm{\epsilon}\cdot\hat{\bm{r}}\big{)}^{2}-\frac{1}{2}t^{2}\epsilon^{2}\Big{]}\;.$ (2.15)
Here, and in the remainder of the paper, we neglect terms of $\mathcal{O}({\epsilon}^{3})$. Consequently, from now on all relevant expressions are implicitly understood to hold up to $\mathcal{O}\big{(}{\epsilon}^{2},\delta(\mathcal{H}/k)^{2}\big{)}$. Using the above results in (2.13), we obtain
@@
Then the full kernel, i.e. standard + combined relativistic and wide-angle corrections, is
@@ 3 Auto- and cross-power spectra
where $P$ is the linear power spectrum of the comoving matter density contrast $\delta$. Here and below, we omit the dependencies on $z$ for brevity. These are the ‘local’ power spectra, i.e. the Fourier power spectra at line-of-sight position $\bm{r}(z)$. The ‘full’ power spectra are then given by an integral over all lines of sight [22],
@@
$\mathcal{K}_{a}(\bm{k}^{\prime},\bm{\epsilon},{\bm{r}})\,\mathcal{K}_{b}^{*}(\bm{k}^{\prime},\bm{\epsilon},{\bm{r}})=\sum_{l,m,n}\,\mathcal{C}_{lmn}^{(ab)}(\bm{k}^{\prime},{\bm{r}})\,\epsilon_{x}^{l}\,\epsilon_{y}^{m}\,\epsilon_{z}^{n}{\color[rgb]{0,0,0}{\quad\mbox{where}\quad l+m+n\leq 2}}\;.$ (3.3)
Here the coefficients $\mathcal{C}_{lmn}^{(ab)}$ are independent of $\bm{\epsilon}$ and the condition on $l,m,n$ excludes terms of $\mathcal{O}(\epsilon^{3})$. Then the power spectrum (3.1) becomes
@@
$=P^{\mathrm{S}}_{ab}+P^{\mathrm{R}}_{ab}+P^{\mathrm{W}}_{ab}+P^{\mathrm{RW}}_{ab}\;.$ (3.10)
@@
Here $\mathcal{L}_{\ell}$ are Legendre polynomials. The detailed expressions for the dominant multipoles $\ell=0,1,2$ are given in Appendix C.
@@ 5 Detecting the relativistic and wide-angle effects
Our aim is to test whether the detection of relativistic and wide-angle effects is feasible in futuristic galaxy surveys. Following [27], we compute the $\chi^{2}$ values between the model $P_{ab}^{\rm S}+P_{ab}^{\rm I}$, which includes the relativistic and/or wide-angle effect I, and the standard model $P_{ab}^{\rm S}$, which excludes the effect I. We consider all the cases of $P_{ab}^{\rm I}$, where I=W, R, RW or W+R+RW. Our aim is to estimate the statistical significance of measuring effect I, against the null hypothesis of the absence of such an effect in the data.
In our case, we use synthetic data generated by taking the theoretical prediction of the full power spectrum (3.10). Therefore the $\chi^{2}$ for effect I in a redshift bin centred at $z_{i}$ is simply
@@
$\mathrm{Var}\big{[}P_{aa}\big{]}=\frac{2}{N_{\bm{k}}}\,\big{|}\tilde{P}_{aa}\big{|}^{2}\,,\quad\mathrm{Var}\big{[}P_{12}\big{]}=\frac{1}{N_{\bm{k}}}\,\left[\,\big{|}\tilde{P}_{12}\big{|}^{2}+\big{|}\tilde{P}_{11}\big{|}\big{|}\tilde{P}_{22}\big{|}\,\right]\;.$ (5.2)
Here
$\tilde{P}_{ab}=P_{ab}+{\color[rgb]{0,0,0}{P_{ab}^{\rm shot}}}\qquad\mbox{and}\qquad P_{ab}=P_{ab}^{\rm S}+P_{ab}^{\rm R+W+RW}\,.$ (5.3)
The shot noise is given by $P_{aa}^{\rm shot}=n_{a}^{-1}$ and $P_{12}^{\rm shot}=0$.
@@
Note that $P_{ab}$ in the variance in (5.1)–(5.2) is given by (3.10) – i.e. it includes all relativistic and wide-angle effects. The number of Fourier modes in a redshift bin centred on $z_{i}$ is
$N_{\bm{k}}{\color[rgb]{0,0,0}{(z_{i})}}=\frac{2\,\pi\,k^{2}\,\Delta k\,\Delta\mu}{k_{\mathrm{f}}{\color[rgb]{0,0,0}{(z_{i})}}^{3}}\;.$ (5.4)
Here the fundamental mode, $k_{\mathrm{f}}{\color[rgb]{0,0,0}{(z_{i})}}$, corresponds to the longest wavelength that can be measured in the redshift bin centred on $z_{i}$. The Fourier modes in each bin satisfy $k_{\rm min}(z_{i})\leq k\leq k_{\rm max}(z_{i})$, where the minimum and maximum modes are given in (5.7) and (5.8).
The $\chi^{2}$ cumulative over redshift bins is
@@ Acknowledgements
SJ is supported by the Stellenbosch University Astrophysics Research Group fund. SG and RM are supported by the South African Radio Astronomy Observatory and National Research Foundation (grant no. 75415). SC acknowledges support from the Italian Ministry of University and Research, PRIN 2022 ‘EXSKALIBUR – Euclid-Cross-SKA: Likelihood Inference Building for Universe Research’, from the Italian Ministry of Foreign Affairs and International Cooperation (grant no. ZA23GR03), and from the European Union – Next Generation EU. We used the Python code Sympy for analytical calculations.
@@ Appendix A Fourier kernels
Recall that throughout the paper we work to $\mathcal{O}\big{[}\epsilon^{2},\delta(\mathcal{H}/k)^{2}\big{]}$.
@@
${}~{}~{}~{}~{}+2(t-\sigma_{a})\Big{[}k_{x}k_{z}\epsilon_{x}+k_{y}k_{z}\epsilon_{y}\Big{]}\Big{\}},$
@@
${}+\frac{3}{2}\Omega_{m}\frac{\mathcal{H}^{2}}{k^{2}}\frac{(1-\mathcal{Q}_{a})}{r\mathcal{H}}\Big{[}2(t-\sigma_{a})\epsilon_{z}+(t-\sigma_{a})^{2}(\epsilon_{x}^{2}+\epsilon_{y}^{2}-2\epsilon_{z}^{2})\Big{]},\qquad$